A random sample of size 100 has a standard deviation of 5.Find

maximum error with 95% confidence.


1
Expert's answer
2022-05-03T13:12:34-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

Maximum error


"E=z_c\times\dfrac{\sigma}{\sqrt{n}}"

"E=1.96\times\dfrac{5}{\sqrt{100}}=0.98"